OpenStudy (anonymous):

If tan x=-4 and x is negative, find sin2x, cos2x and tan2x

Parth (parthkohli):

So many ways to do this...

OpenStudy (anonymous):

well i'm looking for the answer that uses double angle identities

Parth (parthkohli):

\[\sin(2x) = \dfrac{2\tan(x)}{1 + \tan^2(x)}\]\[\cos(2x) = \dfrac{1 - \tan^2(x)}{1 + \tan^2(x)}\]

Parth (parthkohli):

And\[\tan(2x) = \dfrac{2\tan (x)}{1 - \tan^2(x)}\]By dividing those two.

OpenStudy (anonymous):

are those considered identities? I only have identities for \[\cos 2x=\cos ^{2}x-\sin ^{2}x\]

OpenStudy (anonymous):

and \[\sin2x=2sinxcosx\]

Parth (parthkohli):

Yes, these are identities.

OpenStudy (anonymous):

ok thanks. I don't think we've gotten that far yet. Appreciate the help!!

Parth (parthkohli):

OK, I'll try to help you using the identities you already know.

Parth (parthkohli):

You do know the identity for \(\tan(2x)\) though, right?

OpenStudy (anonymous):

yes that one was familiar and in my textbook in this chapter.

Parth (parthkohli):

OK, you can find tan(2x) to begin with then.

OpenStudy (anonymous):

Thats my work. maybe it will help someone else. Thanks again!!

Parth (parthkohli):

OK, that's a valid approach too. Good going.

Parth (parthkohli):

Just a little warning of sorts, though: even though your approach works here, tan(x) = -4 can also mean that sin(x) = 4 and cos(x) = -1. So you should always think twice before proceeding with these things.

OpenStudy (anonymous):

definitely. I was told that SinX was negative in my book and i wrote it down of sorts but it wasn't very clear in the question i typed on this site. I'll be sure to keep an eye out!

Parth (parthkohli):

Oh, that was given? It's a very vital piece of information.

OpenStudy (anonymous):

hehe yea, I wrote "X is negative" when i meant to write "SINX was negative.